# The HLLC-2D method for the computation of two phase flow system ejecta transporting model

### v1

@@ 3 Numerical Scheme
In this section,we will give a numerical scheme to solve the equation (2.1) and (2.2).
@@ 3.1 ALE scheme
For the convenience, recast the equation (2.1) as follows:
where the state vector and flux vectors are
$\mathbf{U}=\left[\begin{array}[]{c}\rho_{g}\\ \rho_{g}v_{g_{x}}\\ \rho_{g}v_{g_{y}}\\ \rho_{g}E_{g}\end{array}\right],\quad\mathbf{F}(\mathbf{U})=\left[\begin{array}[]{c}\rho_{g}v_{g_{x}}\\ P+\rho_{g}v_{g_{x}}^{2}\\ \rho_{g}v_{g_{x}}v_{g_{y}}\\ \rho_{g}E_{g}v_{g_{x}}+Pv_{g_{x}}\end{array}\right],\quad\mathbf{G}(\mathbf{U})=\left[\begin{array}[]{c}\rho_{g}v_{g_{y}}\\ \rho_{g}v_{g_{x}}v_{g_{y}}\\ \rho_{g}v_{g_{y}}^{2}+P\\ \rho_{g}E_{g}v_{g_{y}}+Pv_{g_{y}}\end{array}\right],$ (3.2)
and $\mathbf{S}_{p}$ is the resource term between the flow and particles
$\mathbf{S}_{p}=\left[\begin{array}[]{c}0\\ {F}_{p_{1}}\\ {F}_{p_{2}}\\ {E}_{p}\end{array}\right].$ (3.3)
Based on ALE method, we give the numerical scheme. For the subsequent discretization process, it is advantageous to express equation (2.1) in the form of a moving control volume formulation:
@@
$\mathbf{w}$ represents the moving velocity of the control volume $\Omega$ while $\mathbf{N}$ denotes the unit outward normal direction along the boundary of $\Omega$. When $\mathbf{w}=\mathbf{u}$, the system simplifies into a Lagrangian formulation, and when $\mathbf{w}=\mathbf{0}$, it takes on an Eulerian form.
@@ 3.1.1 Notations on a generic polygonal grid
Every cell $\Omega_{c}$ within the mesh is assigned a distinct index represented by $c$. To refer to an adjacent cell $\Omega_{d}$ that shares a common edge with $\Omega_{c}$, we employ the subscript $d$. The edge where two units intersect is defined as $c\cap d$ or denoted as $k$, as shown in Fig 2. ${F}(c)$ represents the collection of neighboring cells surrounding cell $\Omega_{c}$. and ${Q}(c)$ is the set of all vertices of a cell $\Omega_{c}$. Each vertex within the mesh is allocated a unique index $q$, and $C(q)$ and $K(q)$ are defined as the sets of cells and edges respectively surrounding the vertex $q$, i.e.,
$C(q)=\left\{\Omega_{c}\text{ : cells sharing a vertex }q\right\},$
$K(q)=\{k\text{ : edges sharing a vertex }q\}.$
The physical quantities of flow such as the density $\rho_{g_{c}}$, pressure $P_{g_{c}}$, velocity $\bm{v}_{g_{c}}$, energy $E_{g_{c}},e_{g_{c}}$ are defined in the cell center of $\Omega_{c}$. The grid moving velocity, denoted as $\mathbf{w}_{q}$, is defined on the node $q$. Let $\mathbf{N}_{d}^{c}=\left(\left(n_{x}\right)_{d}^{c},\left(n_{y}\right)_{d}^{c}\right)$ represents the unit normal vector of cell $\Omega_{c}$ along edge $c\cap d$, with the subscript $d$ and the superscript $c$ indicating that the vector direction originates from $\Omega_{c}$ and extends towards $\Omega_{d}$.
@@
Based on the edge’s normal direction between two adjacent cells, we utilize indices $L$ and $R$ to represent the states on the left and right sides of the edge, respectively. For example, $\Omega_{c}$ and $\Omega_{d}$ are the left and right cells of the edge $c\cap d$.
To compute integration of the variable $\mathbf{F}$ across face $\left[M_{a},M_{a+1}\right]$, denote by $\mathbf{F}_{a,a+1}$ the fluxes on the face, $L_{a,a+1}$ is the length of $\left[M_{a},M_{a+1}\right]$, we have
@@
Using the above transformation to discretize the equations (3.4) by utilizing the idea of the Godunov scheme on unstructured mesh. Define cell-centered state variable $\mathbf{U}_{c}$ and resource term $\mathbf{S}_{p_{c}}$, then we have:
@@
where $\left|\Omega_{c}\right|$ represent the volume of cell $\Omega_{c}$. $\mathbf{w}_{d}^{c}$ is the moving velocity which is defined at the center of cell edge $c\cap d$. $L_{d}^{c}$ is the length of the edge $c\cap d$. $\mathcal{T}_{d}^{c}$ and $\left(\mathcal{T}_{d}^{c}\right)^{-1}$ are the rotation matrix and its inverse as follows:
@@
$\mathbf{U}_{t}+\mathbf{F}_{x}=\mathbf{S}_{p}$ (3.8)
@@
where $w=\mathbf{w}_{d}^{c}\cdot\mathbf{N}_{d}^{c}$ is the normal projection of the moving velocity $\mathbf{w}_{d}^{c}$ , $\mathbf{U}^{w}$ and $\mathbf{F}^{w}$ are the projection of the state and flux functions onto the direction $w=x/t$ respectively.
@@ 3.1.2 Flux formula of HLLC-2D on moving mesh
We use the two-dimensional Riemann solver HLLC-2D method proposed in [2, 1] to evaluate the numerical flux. The construction of this methods hinges on two important points. Firstly, it entails a nodal solver that relies on the vertex $q$ with multi-state values $\left\{\mathbf{U}_{c},c\in C(q)\right\}$ assigned to neighbor cells. Then the nodal contact velocity $\mathbf{u}_{q}^{*}$ is determined through this solver. Secondly, it involves the utilization of the HLLC-2D solver, which is an expanded scheme of the classical one-dimensional HLLC Riemann solver. In Riemann problem (3.8), we have the nodal flow velocity $\mathbf{u}_{q}^{*}$, as well as the state vectors $\mathbf{U}_{c}$ and $\mathbf{U}_{d}$, which are on the sides of edge $c\cap d$. Through solving the Riemann problem, we could get the approximate numerical flux $\mathbf{F}_{c}^{d}$ and $\mathbf{F}_{d}^{c}$ and further solving the equation (3.6).
In the following evaluation, assume that the nodal contact velocity $\mathbf{u}_{q}^{*}$ has already been determined, in alignment with the preceding chapter’s description. From the equation (3.8), the initial values of the equation are:
@@
so that the corresponding fluxes are $\mathbf{F}_{L}=\mathbf{F}\left(\mathbf{U}_{L}\right),\mathbf{F}_{R}=\mathbf{F}\left(\mathbf{U}_{R}\right)$.
@@
where $H=L,R$, and the above flux scheme ensure strictly satisfies the R-H conditions.
@@
$\mathbf{F}_{d}^{c}\left(w,\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)=\frac{1}{2}\left(\mathbf{F}_{d,q}^{c}\left(\mathbf{w}_{q}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)+\mathbf{F}_{d,q^{+}}^{c}\left(\mathbf{w}_{q^{+}}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)\right)$ (3.15)
@@ 3.1.3 Nodal Solver
In Subsection 3.1.2, we have assumed that the nodal contact velocity is already calculated and then present a nodal solver that consistently evaluate the $\mathbf{u}_{q}^{*}$ to ensure that the scheme maintains local conservation of mass, momentum and total energy. Since $\mathbf{F}_{d}^{c}\neq\mathbf{F}_{c}^{d}$ may happen, the conservation in cells is not satisfied. Refer to the construction of the solver, we want to ensure the conservation of mass, momentum and total energy enforced at nodes and we could derive this procedure from the global conservation within domain $\Omega$. By maintaining the local nodal conservation including mass, momentum and total energy, contact velocity could be calculated. By using the pressure equation (3.14), we have:
@@
and $v_{k}^{*}$ is the contact velocity in the classical one-dimensional HLLC Riemann solver for edge $k$
@@
Solve the equation (3.16) to calculate the contact velocity $\mathbf{u}_{q}^{*}$ at each vertex $q$. Then we get
@@ 3.1.4 Boundary Condition
In this ALE method, it employs the convensional zero-order extrapolation method to enforce boundary conditions. For velocity-based boundaries, such as solid wall or piston, two distinct cases arise (1) When three boundary nodes are not collinear, we apply a straightforward extrapolation method. (2) In cases where three boundary nodes are collinear, we employ t mirror extrapolation which entails the construction of a symmetrical point for each vertex connecting a boundary node. This method ensures the same scheme (3.19) as used for internal vertex when calculating $\mathbf{u}_{q}^{*}$.
4 Time discretization
In this section, we discretize in time for the system (3.6) that describes the evolution of the variables in the cell $\Omega_{i}$. Suppose that there are $m(i)$ particles in cell $\Omega_{i}$, of which the index is $m=1,\ldots,m(i)$, then we use a second-order Runge-Kutta scheme:
@@ 4 Time discretization
$\overline{\mathbf{U}}_{i}^{n}=\frac{1}{2}({\mathbf{U}}_{i}^{n}+\hat{\mathbf{U}}_{i}^{n+1})$ (4.2)
@@ 4.1 Particle searching
    Suppose that we know the coordinate of particle m at time $t_{n}$, $\bm{u_{m}^{n}}=(x_{m}^{n},y_{m}^{n})$. And the particle m is in the cell $\Omega_{i}$.
@@
    Calculate the equation (4.5) and get the coordinate of particle m, $\bm{u_{m}^{n+1}}=(x_{m}^{n+1},y_{m}^{n+1})$ , at time $t_{n+1}$.
@@
    If not, locate the surrounding cells of $\Omega_{i}$, suppose the index of these cells are $c=1,\dots,n(i)$. Search the particle m in the cell $\Omega_{c}$, $c=1,\dots,n(i)$.
@@
    Update the location of particle m.
Using the above method, we could get the updated location of particle k under the condition that we give a limitation to the time step $\Delta t^{n}$.
@@ 4.2 Time step limitation
Because of using the explicit Euler method, we need to observe some limitations on time step. They are as follows:
@@
    The variation of cell volume is coherent with the vertex motion.
@@
    A CFL-like condition to be a priori satisfied for entropy inequality.
@@
Suppose that we know the physical properties in the cell $\Omega_{i}$ and the geometrical characteristics of the cell $\Omega_{i}$ at the beginning of the time step $t_{n}$. We will compute the variables’ value at time $t_{n+1}$ and time step $\Delta t^{n}$.
@@ 4.2.1 CFL condition
First we derive a CFL condition to ensure a positive entropy production in cell $\Omega_{i}$ during the time step. At time $t_{n}$, for each cell $\Omega_{i}$ we denote by $\lambda_{i}^{n}$ the inner circle radius of the cell. We have
@@
where $C_{E}$ is a strictly positive coefficient and $c_{i}$ is the sound speed in the cell. The coefficient $C_{E}$ is given artificially. According to the numerical experiments ,$C_{E}=0.3$ is a value which could get good numerical results.
@@ 4.2.2 Criterion on the variation of volume
We also give the criterion on the variation of the cell volume. At time $t_{n}$, we can get the contact velocity of vertex, we have:
$\Delta t_{v}=C_{V}\min_{i=1,\ldots,I}\left\{\frac{\left|A_{i}^{n}\right|}{\left|A_{i}^{\prime}\right|}\right\}.\$ (4.7)
where $C_{v}=0.1$ and
@@ 4.2.3 Particle searching limitation
Refer to Subsection 4.1, suppose that particle $m$ is located in $\Omega_{m_{i}}$, we check the cell $\Omega_{m_{i}}$ and its surrounding cells. From time $t_{n}$ to $t_{n+1}$ the particles should be specify the range of motion of the particle. In the subsection 4.2.1 denote by $\lambda_{m_{i}}^{n}$ the inner circle radius of the cell. We have:
$\Delta t_{p}=C_{p}\min_{m}\min_{i=1,\ldots,n(m)}\frac{\lambda_{m_{i}}^{n}}{|\bm{v_{p_{m}}^{n}}|}.\$ (4.9)
where $n(m)$ is the number of cells that is containing and surrounding the particle. In numerical simulation we take $C_{p}=0.5$.
@@
$\Delta t^{n}=\min\left(\Delta t_{e},\Delta t_{v},\Delta t_{p},1.01*\Delta t^{n-1}\right),\$ (4.10)
4.3 Algorithm for discretization

### v2

@@ 3 Numerical Scheme
Then we give a numerical scheme to solve the equation (2.1) and (2.3).
@@ 3.1 ALE scheme
For the subsequent discretization process, it is convenient to express equation (2.1) in the form of a moving control volume formulation:
@@
$\mathbf{w}$ denotes the moving velocity of the control volume $\Omega$ while $\mathbf{N}$ denotes the unit outward normal direction along the boundary of $\Omega$. When $\mathbf{w}=\mathbf{u}$, the system simplifies into a Lagrangian formulation, and when $\mathbf{w}=\mathbf{0}$, it takes on an Eulerian form.
@@ 3.1.1 Notations on a generic polygonal grid
Every cell $\Omega_{c}$ within the mesh is assigned a distinct index represented by $c$. To refer to an adjacent cell $\Omega_{d}$ that shares a common edge with $\Omega_{c}$, we employ the subscript $d$. The edge where two units intersect is defined as $c\cap d$ or denoted as $k$, as shown in Fig 1. ${F}(c)$ denotes the set of adjacent cells surrounding the cell $\Omega_{c}$. ${Q}(c)$ is the set of all vertices of a cell $\Omega_{c}$. Each vertex within the mesh is allocated the index $q$, and $C(q)$ and $K(q)$ are defined as the sets of cells and edges respectively surrounding the vertex $q$, i.e.,
$C(q)=\left\{\Omega_{c}\text{ : cells surrounding the vertex }q\right\},$
$K(q)=\{k\text{ : edges surrounding the vertex }q\}.$
The physical quantities of fluid for example the density $\rho_{g_{c}}$, pressure $P_{g_{c}}$, velocity $\bm{v}_{g_{c}}$, energy $E_{g_{c}},e_{g_{c}}$ are defined in the center of $\Omega_{c}$. The grid moving velocity, denoted as $\mathbf{w}_{q}$, is defined at the node $q$. Let $\mathbf{N}_{d}^{c}=\left(\left(n_{x}\right)_{d}^{c},\left(n_{y}\right)_{d}^{c}\right)$ represents the unit normal vector of cell $\Omega_{c}$ along edge $c\cap d$, with the subscript $d$ and the superscript $c$ indicating that the vector direction originates from $\Omega_{c}$ and extends towards $\Omega_{d}$.
@@
Based on the edge’s normal direction between two adjacent cells, we utilize indices $L$ and $R$ to represent the states on the left and right sides of an edge, respectively. For instance, $\mathbf{U}_{L}$ and $\mathbf{U}_{R}$ are state vectors in the left and right cells of the edge $c\cap d$.
To compute integration of the variable $\mathbf{F}$ across edge $\left[M_{a},M_{a+1}\right]$, denote by $\mathbf{F}_{a,a+1}$ the fluxes on the face, $L_{a,a+1}$ is the length of $\left[M_{a},M_{a+1}\right]$, we have
@@
Using the above transformation, we discretize the equations (3.1) by utilizing the Godunov scheme on an unstructured mesh. Define cell-centered state variable $\mathbf{U}_{c}$ and resource term $\mathbf{S}_{p_{c}}$, then we have:
@@
where $\left|\Omega_{c}\right|$ represent the volume of cell $\Omega_{c}$. $\mathbf{w}_{d}^{c}$ is the moving velocity defined at the center of cell edge $c\cap d$. $L_{d}^{c}$ is the length of the edge $c\cap d$. $\mathcal{T}_{d}^{c}$ and $\left(\mathcal{T}_{d}^{c}\right)^{-1}$ are the rotation matrix and its inverse as follows:
@@
$\mathbf{U}_{t}+\mathbf{F}_{x}=\mathbf{S}_{p},$ (3.5)
@@
where $w=\mathbf{w}_{d}^{c}\cdot\mathbf{N}_{d}^{c}$ is the normal projection of the moving velocity $\mathbf{w}_{d}^{c}$ , $\mathbf{U}^{w}$ and $\mathbf{F}^{w}$ are the projection of the state and flux functions onto the direction $x/t=w$ respectively.
@@ 3.1.2 Flux formula of HLLC-2D on moving mesh
We use the two-dimensional Riemann solver HLLC-2D method proposed in [34] to evaluate the numerical flux. The construction of this method hinges on two important points. Firstly, it involves a nodal solver that relies on the vertex $q$ with state vectors $\left\{\mathbf{U}_{c},c\in C(q)\right\}$ assigned to neighbor cells. Then the nodal contact velocity $\mathbf{u}_{q}^{*}$ is computed through this solver. Secondly, it involves the utilization of the HLLC-2D solver, an expanded scheme of the classical one-dimensional HLLC Riemann solver. In Riemann problem (3.5), we have the nodal contact velocity $\mathbf{u}_{q}^{*}$, as well as the state vectors $\mathbf{U}_{c}$ and $\mathbf{U}_{d}$, which are on the sides of edge $c\cap d$. Through solving the Riemann problem, we could get the approximate numerical flux $\mathbf{F}_{c}^{d}$ and $\mathbf{F}_{d}^{c}$ and further solving the equation (3.3).
In the following evaluation, assume that the nodal contact velocity $\mathbf{u}_{q}^{*}$ has already been calculated, in alignment with the preceding chapter’s description. From the equation (3.5), the initial values of the equation are:
@@
and the fluxes are $\mathbf{F}_{L}=\mathbf{F}\left(\mathbf{U}_{L}\right),\mathbf{F}_{R}=\mathbf{F}\left(\mathbf{U}_{R}\right)$.
@@
where $H=L,R$, and the above flux scheme ensure strictly satisfies the Rankine-Hugoniot conditions.
@@
$\mathbf{F}_{d}^{c}\left(w,\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)=\frac{1}{2}\left(\mathbf{F}_{d,q}^{c}\left(\mathbf{w}_{q}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)+\mathbf{F}_{d,q^{+}}^{c}\left(\mathbf{w}_{q^{+}}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)\right).$ (3.12)
@@ 3.1.3 Nodal Solver
In Subsection 3.1.2, we have assumed that the nodal contact velocity is already calculated. And since $\mathbf{F}_{d}^{c}\neq\mathbf{F}_{c}^{d}$ may happen, the conservation in cells is not satisfied. Therefore we introduce a nodal solver that consistently evaluates the $\mathbf{u}_{q}^{*}$ , ensuring the scheme preserves local conservation of mass, momentum and energy. This procedure can be derived from the global conservation within domain $\Omega$. By maintaining the local nodal conservation including mass, momentum and total energy, contact velocity could be calculated. By using the pressure equation (3.11), we have:
@@
and $v_{k}^{*}$ represents the contact velocity in the classical one-dimensional HLLC Riemann solver for edge $k$
@@
Solve the equation (3.13) to calculate the contact velocity $\mathbf{u}_{q}^{*}$ at each vertex and we have
@@ 3.1.4 Boundary Condition
In this ALE method, the conventional zero-order extrapolation method is employed to enforce boundary conditions. For velocity-based boundaries, there are two different cases:
  • (1)
    When the three boundary nodes are not collinear, a straightforward extrapolation method is implemented.
  • (2)
    When the three boundary nodes are collinear, we employ the mirror extrapolation method, which involves constructing a symmetrical point for each vertex connected to a boundary node.
The above method ensures the same scheme (3.16) as used for internal vertex when calculating $\mathbf{u}_{q}^{*}$.
4 Time discretization and particle searching
We discretize in time for the system (3.3) that describes the evolution of the variables in the cell $\Omega_{i}$. Suppose that there are $m(\Omega_{i})$ particles in cell $\Omega_{i}$, of which the index is $m=1,\ldots,m(\Omega_{i})$, then we use the second-order Runge-Kutta method to obtain the following fully discretized system:
@@ 4 Time discretization and particle searching
$\overline{\mathbf{U}}_{i}^{n}=\frac{1}{2}({\mathbf{U}}_{i}^{n}+\hat{\mathbf{U}}_{i}^{n+1}),$ (4.2)
@@ 4.1 Particle searching
    Suppose that we know the coordinates of particles in the cell $\Omega_{i}$ at time $t_{n}$, namely $\bm{u_{m}^{n}}$, $m=1,\ldots,m(\Omega_{i})$.
@@
    Calculate the equation (4.5) and get the coordinate of particle m, $\bm{u_{m}^{n+1}}$ , at time $t_{n+1}$.
@@
    If not, search the particle m in the cell $\Omega_{c}$ which is the surrounding cells of $\Omega_{i}$, $c=1,\dots,n(i)$.
@@
    Update the position of particle m.
Using the above method, we could get the updated position of particle k under the condition that we give a limitation to the time step $\Delta t^{n}$. In figure 2 we give the schematic diagrams of particle searching on structured and unstructured mesh respectively.
@@ 4.2 Time step limitation
Because of using the explicit method, we need to observe some limitations on time step. They are as follows:
@@
    A CFL condition to be a priori satisfied for entropy inequality.
@@
    The variation of cell volume corresponds coherently with the motion of the vertices.
@@
Assume that the physical properties and geometric characteristics of the cell $\Omega_{i}$ are known at the time $t_{n}$. We will compute the variables’ value at time $t_{n+1}$.
@@ 4.2.1 CFL condition
First we derive a CFL condition to ensure positive entropy production within cell $\Omega_{i}$ during the computation. At time $t_{n}$, we denote by $\lambda_{i}^{n}$ the inner circle radius of the cell $\Omega_{i}$. Then
@@
where $C_{E}$ is a strictly positive coefficient and $c_{i}$ is the sound speed of the cell $\Omega_{i}$. According to the numerical experiments, the value of $C_{E}$ between $0.3~{}0.4$ usually could get a good numerical results
@@ 4.2.2 Criterion on the variation of volume
We also establish a criterion for the variation of cell volume. At time $t_{n}$, we can get the contact velocity of vertex, we have:
$\Delta t_{v}=C_{V}\min_{i=1,\ldots,I}\left\{\frac{\left|A_{i}^{n}\right|}{\left|A_{i}^{\prime}\right|}\right\},\$ (4.7)
where $C_{v}$ can be set $0.1$ and
@@ 4.2.3 Particle searching limitation
Refer to Subsection 4.1, suppose that particle $m$ is located in $\Omega_{i}$, we check the cell $\Omega_{i}$ and its surrounding cells $\Omega_{c}$, $c=1,\dots,n(i)$. From time $t_{n}$ to $t_{n+1}$ the particles should be specify the range of motion of the particle. In the subsection 4.2.1 denote by $\lambda_{k}^{n}$ the inner circle radius of the cell $\Omega_{i}$ and the cell $\Omega_{c}$. We have:
$\Delta t_{p}=C_{p}\min_{m}\min_{k=1,\ldots,n(i)+1}\frac{\lambda_{k}^{n}}{|\bm{v_{p_{m}}^{n}}|},\$ (4.9)
where $n(m)$ is the number of cells containing and surrounding the particle. In numerical simulation we take $C_{p}=0.5$.
@@
$\Delta t^{n}=\min\left(\Delta t_{e},\Delta t_{v},\Delta t_{p},1.01*\Delta t^{n-1}\right).\$ (4.10)
4.3 Algorithm for full discretization system